OpenStudy (anonymous):

Write the following power series as a function d) \(\sum_{n=1}^{\infty}\frac{1}{n}x^{n+3}\)

OpenStudy (blockcolder):

\[\sum_{n=1}^{\infty}\frac{1}{n}x^{n+3}=x^3\sum_{n=1}^{\infty}\frac{1}{n}x^{n}=x^3\sum_{n=1}^{\infty}\left (\int x^{n-1}\ dx\right )\]

OpenStudy (anonymous):

is it end solution? or it needs to be add something more?

OpenStudy (blockcolder):

No, it's not the end yet. \[x^3\sum_{n=1}^{\infty}\left (\int x^{n-1}\ dx\right )=x^3 \int \left ( \sum_{n=1}^\infty x^{n-1}\right )dx\] Can you see what to do now?

OpenStudy (anonymous):

no not really, but you can go on with your solution, i try to figure it out

OpenStudy (blockcolder):

\[\sum_{n=1}^\infty x^{n-1}= \frac{1}{x^2}\sum_{n=0}^\infty x^n=\frac{1}{x^2(1-x)}\] Now integrate this, and multiply by x^3 and that's the answer.

OpenStudy (anonymous):

ok i will try but i am not sure if can if you can bring it to end it will be much appreciate, i am noob at mathematics and must give homework soon

OpenStudy (anonymous):

:(

OpenStudy (blockcolder):

\[\frac{1}{x^2(1-x)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{1-x}\\ 1=A(x)(1-x)+B(1-x)+Cx^2\\ 1=Ax-Ax^2+B-Bx+Cx^2\\ 1=(C-A)x^2+(A-B)x+B \\ A=B=C=1\\ \int \frac{1}{x^2(1-x)}\ dx= \int \frac{1}{x}+\frac{1}{x^2}+\frac{1}{1-x}\ dx\\ =\ln\left (\frac{x}{1-x} \right )-\frac{1}{x}\\ \sum_{n=1}^{\infty}\frac{1}{n}x^{n+3}=x^3 \ln\left (\frac{x}{1-x} \right )-x^2\]

OpenStudy (anonymous):

now this is end solutioin right?

OpenStudy (blockcolder):

It's the end.

OpenStudy (anonymous):

ok thank you very very much